mathslover (mathslover):

Solve for \(x\in R\) : \[ e^{\tan x} + \sec x = \log ( \cot x )+ \csc x \]

mathslover (mathslover):

I have no idea from where to start this. \[\large{ e^{\tan x} - \log \left( \cfrac{1}{\tan x} \right) = \csc x - \sec x }\]

mathslover (mathslover):

\(\color{blue}{\text{Originally Posted by}}\) @mathslover \[ \large{ e^{\tan x} - \log \left( \tan (x) \right) ^{-1} = \csc x - \sec x \\ e^{\tan x} + \log \left(\tan x \right) = \csc x - \sec x } \] is this okay? \(\color{blue}{\text{End of Quote}}\)

mathslover (mathslover):

@experimentX @BSwan and @Miracrown

OpenStudy (experimentx):

do you want exact solution or approximate solution?

mathslover (mathslover):

Not sure, how will you do to solve this for approximate solution. If you think that finding approximate solution will be okay then I would love to see it. [I saw this question on Facebook, and posted it here]

OpenStudy (experimentx):

there is a well known method on numerical analysis. http://en.wikipedia.org/wiki/Newton's_method

OpenStudy (experimentx):

wolf can't find it ... "standard computation time exceeded" http://www.wolframalpha.com/input/?i=solve++e%5Etan%28x%29+%2B+sec%28x%29+%3D+log%28cot%28x%29%29+%2B+csc%28x%29

OpenStudy (anonymous):

try to find a,b s,t f(a)=-ve f(b)=+ve then every thing is well be good using any approximation

mathslover (mathslover):

Well, someone commented there that is it : \(\large e^{\tan x + \sec x} = \log (\cot x + \csc x) \) I'm not sure whether this is the question or the above one. I'm sorry.

mathslover (mathslover):

I never studied Newton's Method. I will have to give some time to it though. But, @experimentX and @BSwan , do you both think that this will be the best method to solve such questions?

OpenStudy (experimentx):

the newton's method is numerical method, if you can avoid it, avoid it.

OpenStudy (anonymous):

there is no best way :) if u can solve exactly then its the best way

mathslover (mathslover):

Okay. Well, Let us just suppose that the question was : \[\large e^{\tan x + \sec x} = \log (\cot x + \csc x ) \] Any particular ideas, on how to solve this?

OpenStudy (anonymous):

im thinking :O i would chick domain/range for both

OpenStudy (experimentx):

simpification yields e^((sin(x) + 1)/cos(x)) = log((cos(x) + 1)/sin(x)) e^(f(x)) = log(f(pi/2-x))

mathslover (mathslover):

Oh... nice

OpenStudy (experimentx):

the equality occurs when f(x) = -infinity and f(pi/2 - x) = 1

mathslover (mathslover):

I have a question, what level do you think the question belongs to?

OpenStudy (experimentx):

i honestly don't know ... mathematica can't find it's numerical solution.

mathslover (mathslover):

That one is for the assumed quest. right? (second quest.)

OpenStudy (experimentx):

huh?

mathslover (mathslover):

\(e^{\tan x + \sec x} = \log (\cot x + \csc x) \) The statement you gave above was for this question? ^

OpenStudy (experimentx):

yep

mathslover (mathslover):

Okay. Though, I'm not getting it completely , but I will again refer to Newton's Method as this seems to be much higher for me, and may be, it would help me in future if I learn Newton's Method. Thanks @experimentX .

mathslover (mathslover):

Oh...Okay! :(

mathslover (mathslover):

Yeah... seems to be like 0.49 or 0.48.

OpenStudy (experimentx):

hmm ... don't be deceived, it can be irrational as well.

mathslover (mathslover):

Okay! I think, I am getting now..

OpenStudy (experimentx):

ok ok ... i don't feel like doing it right now lol. I think numerical method is the best to do it right now. consider let the expression be f(x) = g(x) then let y = f(x) - g(x) and use newton's method. and best of luck :) gotta go ... see you later.

mathslover (mathslover):

Sure. Thanks!